OpenStudy (molly98):

A set of 14 data is: 6, 8, 7, 7, 5, 7, 6, 8, 6, 9, 7, p, q The mean and mode of the set are both 7. Find and q.

OpenStudy (molly98):

anyone got any ideas?

OpenStudy (will.h):

what is the question what's find and q ?

OpenStudy (molly98):

what?

OpenStudy (molly98):

the question is to find p and q.

OpenStudy (sshayer):

6+8+7+7+5+7+6+8+6+9+7+p+q=14*7=98 there are only 13 numbers ,please check

jhonyy9 (jhonyy9):

6, 8, 7, 7, 5, 7, 6, 8, 6, 9, 7, p, q 6+8=14 8+7=15 7+7=14 7+5=12 5+7=12 7+6=13 6+8=14 8+6=14 6+9=15 9+7=16 7+9=16 9+8=17 8+10=18 so p=9 and q=8 than my logic way is right